find the percentage of water (to the nearest tenth) in a crystal of cuso4 x 5h2o
1
Expert's answer
2013-02-26T10:19:30-0500
Percentage of water is next:

w of H2O = 5 * Mw of H2O / Mw CuSO4*5H2O

Mw is molecular mass:

Mw(H2O) = 18 = 90

Mw(CuSO4*5H2O) = 160 + 90 = 250

w of H2O = 90/250*100% = 36 %